Examine the following book-value balance sheet for University Products Inc. The preferred stock currently sells for $15 per share and pays a dividend of $3 a share. The common stock sells for $20 per share and has a beta of 0.7. There are 2 million common shares outstanding. The market risk premium is 12%, the risk-free rate is 8%, and the firm’s tax rate is 21%.

BOOK-VALUE BALANCE SHEET
(Figures in $ millions)
Assets Liabilities and Net Worth
Cash and short-term securities $ 1.0 Bonds, coupon = 6%, paid annually
(maturity = 10 years, current yield to maturity = 8%)
$ 15.0
Accounts receivable 4.0 Preferred stock (par value $20 per share) 3.0
Inventories 8.0 Common stock (par value $0.10) 0.2
Plant and equipment 24.0 Additional paid-in stockholders’ equity 10.8
   Retained earnings 8.0
Total $ 37.0 Total $ 37.0

a. What is the market debt-to-value ratio of the firm? I know the answer for this part which is 23.51%

b. What is University’s WACC?

(For all the requirements, do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)
Answer #1
Bonds (A) 15
Preferred Stock (B) 3
Common stock 0.2
Additional paid-in stockholders’ equity 10.8
Retained earnings 8
Shareholders Equity Capital (C) 19
Total Capital (A+B+C) 37
Source of capital Formula for cost of capital Cost of Capital Capital amount ($ millions) % weightage weightage * cost
Bonds Cost of debt (Post tax) = Pre-tax interest rate * (1- tax rate) 6%*(1-0.21)=4.74% 15 15/37 = 40.5405% 40.5405% * 4.74% = 1.9216%
Preferred stock Cost of preferred stock = Dividend per share / Par value per share 3/20 = 15% 3 3/37 = 8.1081% 8.1081% * 15% = 1.2162%
Equity stock Cost of equity = Risk free rate + Beta * (market premium) 8%+(0.7*12%) = 16.4% 19 19/37 = 51.3514% 51.3514% * 16.4% = 8.4216%
WACC = 1.9216% + 1.2162% + 8.4216%
WACC = 11.56%